# Can Code-Switched Texts Activate a Knowledge Switch in LLMs? A Case Study on English-Korean Code-Switching

### v1

@@ Abstract
Code-switching (CS), a phenomenon where multilingual speakers alternate between languages in a discourse, can convey subtle cultural and linguistic nuances that can be otherwise lost in translation. Recent state-of-the-art multilingual large language models (LLMs) demonstrate excellent multilingual abilities in various aspects including understanding CS, but the power of CS in eliciting language-specific knowledge is yet to be discovered. Therefore, we investigate the effectiveness of code-switching on a wide range of multilingual LLMs in terms of knowledge activation, or the act of identifying and leveraging knowledge for reasoning. To facilitate the research, we first present EnKoQA, a synthetic English-Korean CS question-answering dataset. We provide a comprehensive analysis on a variety of multilingual LLMs by subdividing activation process into knowledge identification and knowledge leveraging. Our experiments demonstrate that compared to English text, CS can faithfully activate knowledge inside LLMs, especially on language-specific domains. In addition, the performance gap between CS and English is larger in models that show excellent monolingual abilities, suggesting that there exists a correlation with CS and Korean proficiency.
@@ 1 Introduction
Code-switching (CS), or the practice of alternating between two or more languages or language varieties within an utterance, is a common phenomenon in multilingual societies. There are multiple motivations for code-switching, to compensate for lack of language proficiency, to emphasize certain emotions or points, or for group identity (Heredia and Altarriba, 2001; Doğruöz et al., 2021).
@@
In particular, code-switching is an effective tool to embedded cultural meanings for bilinguals. Expressing certain concepts in original language can convey subtle cultural and linguistic nuances that can be lost in translation, and knowledge related to certain language are likely to be more memorized in its own language than in foreign languages. As shown in Figure 1, when a human English-Korean bilingual is given a question that is closely related to Korean culture, a question in English and Korean code-switching is more capable of recalling knowledge about “몽유도원도”111A landscape painting by An Gyeon in the early Joseon Dynasty requested by Prince Anpyeong, after his dream about Shangri-la. The painting is drawn on silk with ink., because the concept is more familiar in Korean than in English.
Despite recent attempts to evaluate large language models (LLMs) on CS scenarios, the effectiveness of CS in terms of eliciting language-specific knowledge from LLMs has not been explored. There have been continuous, if not abundant, researches on code-switching in the field of computational linguistics (Aguilar et al., 2020; Rizvi et al., 2021). Recently, after the emergence of LLMs with impressive multilingual abilities, a line of work have discovered LLMs’ abilities in CS (Huzaifah et al., 2024; Yong et al., 2023; Zhang et al., 2023a). However, the focus of such works are only limited to understanding and generating CS of LLMs.
@@
In light of this, we investigate whether code-switching can effectively activate embedded knowledge in LLMs. By knowledge activation, we refer to the overall process of identifying what knowledge is required, and applying knowledge to answer the question. Therefore, we ask ourselves the following research question: Can code-switched texts activate language-specific knowledge, or turn on a “knowledge switch” in LLMs?
@@
To answer the question, we subdivide knowledge activation process into two tasks: (1) In Knowledge Identification task, we investigate if querying LLMs in CS and English yield different knowledge from its encoded memory. Specifically, we evaluate the quality of knowledge from different language settings in terms of faithfulness and helpfulness. (2) In Knowledge Leveraging task, we observe if LLMs can faithfully ground on identified knowledge in different language settings. We evaluate LLMs’ accuracy on question-answering (QA) when given with knowledge.
@@
Meanwhile, a crucial challenge when it comes to code-switching is the data scarcity. There is a limited number of CS datasets, let alone culture-focused data (Doğruöz et al., 2021). Since CS often happens in conversations, CS data are not easily available and the quality is not ensured. Manually collecting CS text is expensive and tricky. To address the shortage of CS data, efforts have been made to synthetically generate data based on linguistic theories (Pratapa et al., 2018; Rizvi et al., 2021; Salaam et al., 2022). However, these works rely on syntactic parsers and part-of-speech taggers that support a limited amount of languages, and the quality of text are highly dependent on the performances of those tools.
Therefore, we first construct EnKoQA, an English-Korean code-switching dataset that follows the Matrix Language Frame Model (Myers-Scotton, 1997). We automatically translate Korean QA datasets that encompass various aspects of Korea (Kim et al., 2024b; Son et al., 2024) into English and synthesize them to construct English-Korean code-switched questions.
@@
We conduct experiments with EnKoQA and provide extensive analysis on a wide range of multilingual LLMs. The experimental results reveal that CS is able to faithfully activate language-specific knowledge that are encoded in multilingual LLMs compared to English; this tendency was more prominent on domains that involve language-specific knowledge. We also discover that proficiency on each language significantly influence knowledge activation performance.
@@
To the best of our knowledge, this work is the first to analyze the effectiveness of code-switching on knowledge activation to LLMs by introducing two tasks.
We propose a qualified English-Korean code-switching QA dataset that is constructed upon two Korean-centric datasets, and conduct extensive experiments on various families of multilingual LLMs.
Experimental results on extensive LLMs indicate that code-switching has advantages in knowledge activation especially on language-specific domains, suggesting the potential of code-switching text as a tool for conveying cultural nuances.
@@ 2.1 Code-Switching Theories
Many linguistic theories attempt to explain the grammatical construction of code-switched text, such as Equivalence Constraint theory and Free Morpheme Constraint theory proposed by Poplack (1980). Equivalence Constraint (EC) theory suggests that code-switching occurs at points in a sentence where the structures of both languages are grammatically compatible. Free Morpheme Constraint (FMC) theory suggests that code-switching cannot occur between a bound morpheme and a lexical base. (e.g., “He is look-ando for a book.” is a wrong code-switch.)
@@ 2.4 Code-Switched Data Synthesis
| GPT-4o | CS | 91.53 | 78.41 | 69.04 | 74.79 | 55.86 | 90.48 | 95.12 | 63.7 | 85.14 | 78.23 |
| EN | 89.83 | 75.00 | 66.19 | 61.97 | 52.64 | 84.52 | 95.12 | 60.4 | 74.32 | 73.33 | |
@@
| EN | 76.27 | 46.02 | 49.47 | 38.46 | 42.30 | 69.05 | 73.17 | 52.15 | 63.51 | 56.71 |
3 EnKoQA: English-Korean Code-Switching QA Dataset
In this section, we introduce EnKoQA, a qualified English-Korean code-switching dataset that is built upon two existing Korean QA datasets. We first discuss the details of data construction (§ 3.1), and conduct experiments to compare performances of LLMs on CS and English questions (§ 3.2, 3.3).
3.1 Dataset Construction
Data Sources.
We leverage two multiple-choice Korean-centric question-answering datasets that encompass various aspects of Korean language and culture. CLIcK (Kim et al., 2024b) consists of 1,995 multiple-choice QA pairs, classified in two main categories (Culture, Language) and 11 sub-categories. CLIcK is sourced from various official Korean exams and textbooks, e.g., College Scholastic Ability Test of Korea (CSAT). In this work, we only utilize data of eight sub-categories from Korean Culture category as our work aims to evaluate the effect of CS on activating Korean-specific knowledge. HAE-RAE (Son et al., 2024) is a Korean benchmark dataset originally crafted to capture cultural and contextual nuances inherent to the Korean language. HAE-RAE is sourced from text on the internet and official exams as well. We use 1,027 multiple-choice QA pairs regarding Korean culture. We combine two datasets and merge common categories (i.e., Society, Geography, and Law), resulting in 2,372 QA pairs in nine categories: Popular, Economy, Politics, Tradition, General Knowledge, Society, Geography, History, and Law. More details of original datasets are provided in Appendix A.1.
@@ 3.1 Dataset Construction
As most LLMs are trained on English-dominant corpora, we regard the multilingual LLM as a bilingual whose matrix language is English but also competent in Korean. To generate code-switched text that follows the MLF model, we need parallel data in Korean and English to extract semantically important words or phrases from Korean text and embed into English text. In order to do so, we first automatically translate all Korean query-choices pairs into English using gpt-3.5-turbo. The model is instructed to translate the following {query} and {choices} to English with an one-shot demonstration in a desired output format. In this paper, we henceforth refer to query-choices pair as question.
@@
Now that we obtain parallel data in both languages, the next step is to embed Korean content morphemes into English sentence. However, as code-switching mostly happens spontaneously, there does not exist a certain formula for mixing two languages. Moreover, replacing every content word with its Korean equivalent may seem rather artificial. To address this, we generate candidate text across four varying proportions of Korean content words: 30%, 50%, 70%, and 90%. Specifically, given a question in both languages and a specified proportion, gpt-3.5-turbo identifies content words from the Korean question and integrates them into the English question according to the specified proportion. To collect contexts of various semantic importance, we employ two prompts that define “content word” differently; one defines content words as nouns or noun phrases, while the other identifies them as semantically important elements within the context. Through this process, we collect total eight code-switched candidates per question. Detailed prompts for each step are provided in Table 4.
Quality Control.
To collect faithful samples, we employ two qualified annotators who are fluent in both Korean and English, and have sufficient background knowledge about South Korea. Given eight code-switched question candidates and a guideline, the annotators examine carefully and choose a single candidate that faithfully follows English sentence structure, and code-switch naturally. We provide a quality control guideline in Appendix A.3, and a sample question in Table 3.
@@ 3.2 Experimental Settings
Implementation Details.
We evaluate performances of LLMs on English and CS questions. For English questions, we use English-translated question from Section 3.1. We simply ask the model to read the following {question} and choose the most appropriate answer, with an one-shot demonstration in a desired output format. The full-length prompts are provided in Table 5.
@@ 3.3 Results
As shown in Table 1, the performance of CS significantly outperforms that of English across many LLMs. The gap is more prominent in models that are recognized for excellent multilingual abilities, such as GPT-4o, Claude 3.5 and Solar. Notably, in the case of Solar, while its performance in English is mediocre, it surpasses Llama3 70B in the code-switching setting.
@@
It is also worth noting that domains where the gap between performance of CS and English is large, such as History and Tradition, are language-sensitive domains. Even for Llama3 and Gemma2 models which do not perform well on CS questions, show higher scores on CS for such domains. On the other hand, domains that are relatively general (e.g., Society, General), and domains that require expert-level knowledge (e.g., Politics, Law), show little or no increase. This finding indicates that asking LLMs in CS is much more effective when it comes to activating Korean-centric knowledge.
Comparison with Korean.
Intuitively, a model’s monolingual capability will affect code-switching on a significant scale. Similarly, we observe that Solar, which was instruction-tuned on Korean datasets (Kim et al., 2024a), outperforms Llama3 70B in CS questions. Thus, we additionally examine the correlation of performances between CS, English, and Korean questions. In Figure 2, models that are located above the diagonal line perform better in Korean than English, and vice versa for the models below the line. The plot shows very consistent trend on all models. It is observable that models that perform better in CS than in English (red markers) are located dominantly above the diagonal line. Meanwhile, models that show low scores on CS, such as Llama3 70B, Gemma2 27B, also perform worse in Korean. In other words, models that perform better on CS also performs well in Korean setting, indicating that performance on CS setting is closely related to Korean proficiency.
@@ 4 Can Code-Switched Questions Activate a “Knowledge Switch” in LLMs?
From Section 3.3, we observe that most LLMs are able to answer correctly to questions in CS than in English. To further investigate on the effectiveness of CS in activating language-specific knowledge, we formulate two tasks: Knowledge Identification and Knowledge Leveraging. An illustration of the two tasks is provided in Figure 3.
4.1 Task Description
Knowledge Identification.
When a human English-Korean bilingual is given a question about Korean culture, they will first try to identify what specific knowledge is required to answer the question, and then apply the knowledge to find the correct answer. Depending on which language the question is written in, the quantity and quality of the knowledge may vary, as described in Figure 1. Language-specific knowledge is likely to be encoded much abundantly in its own language, so reading the question in CS will allow more effective knowledge activation than in English. In this sense, Knowledge Identification task evaluates LLMs’ ability to identify what knowledge is required for answering the question.
Specifically, the LLM is given a question and asked to generate a list of knowledge that are necessary for answering the question. For a qualitative analysis on knowledge identification, we conduct human evaluation on faithfulness and helpfulness of each knowledge list. We will further discuss evaluation details in Section 4.2.
Knowledge Leveraging.
@@ 4.1 Task Description
4.2 Experimental Settings
Evaluation for Knowledge Identification.
In the Knowledge Identification task, we evaluate the quality of a knowledge list based on two criteria: Faithfulness evaluates whether the generated knowledge is factually correct and the model does not output hallucination. Helpfulness evaluates whether the knowledge is helpful for answering the question correctly. A helpful knowledge list should be relevant to the question, and contains the entire knowledge that are necessary for answering the question.
We employ four evaluators who are fluent in both Korean and English and knowledgeable about Korean culture and society. To effectively compare the performance of CS and English, we refer to results from Section 3.3 and choose two domains where CS performance is higher (i.e., History, Tradition), and two domains that have minimum difference (i.e., General, Law). Moreover, we select four models (i.e., GPT-4o, Solar, Gemma2 27B, Gemma2 9B) with different performances and sizes.
We evaluate faithfulness and helpfulness of knowledge on CS and English settings individually, and conduct a pairwise comparison to evaluate which is better. Specifically, we sample 10 questions from each domain and model resulting in total 160 questions. For a single evaluation process, the evaluator is given a question and knowledge list that are created from CS and English settings, and asked to carry out five evaluations, that is, faithfulness/helpfulness of two knowledge lists and pairwise comparison. We provide knowledge lists in a random order. In score-based evaluation, the evaluator is asked to rate each knowledge list on a scale from 1 to 3, where 1 is not faithful/helpful at all, and 3 is significantly faithful/helpful. In pairwise evaluation, we ask the evaluator to select a list that is overall more effective for answering the question; if identical, the evaluator can answer Tie.

### v2

@@ Abstract
Recent large language models (LLMs) demonstrate multilingual abilities, yet they are English-centric due to dominance of English in training corpora. The limited resource for low-resource languages remains a crucial challenge. Code-switching (CS), a phenomenon where multilingual speakers alternate between languages in a discourse, can convey subtle cultural and linguistic nuances that can be otherwise lost in translation and elicits language-specific knowledge in human communications. In light of this, we investigate whether code-switching can activate, or identify and leverage knowledge for reasoning when LLMs solve low-resource language tasks. To facilitate the research, we first present EnKoQA, a synthetic English-Korean CS question-answering dataset. We provide comprehensive analysis on a variety of multilingual LLMs by subdividing activation process into knowledge identification and knowledge leveraging. Our results demonstrate that compared to English text, CS can faithfully activate knowledge inside LLMs especially on language-specific domains, suggesting the potential of code-switching on low-resource language tasks.
@@ 1 Introduction
Large language models (LLMs) have continuously evolved through time to exhibit advanced multilingual capabilities, enabled by training on massive datasets that include text in many different languages. However, these sources are typically skewed toward English, creating an inconsistent performance across different languages (Chen et al., 2024; Zhang et al., 2024). The limited availability for real-world user queries in low-resource languages remains a crucial challenge for achieving robust multilingual models. Prior works attempt to mitigate this issue through machine translation (Artetxe et al., 2023; Bareiß et al., 2024), but crucial semantic nuances may be lost in translation, and machine translation errors are inevitable.
@@
In human multilingual societies, code-switching (CS), or the practice of alternating between two or more languages within an utterance, is used to fill in lack of language proficiency, to emphasize certain emotions or points, or for group identity (Heredia and Altarriba, 2001). Moreover, code-switching functions as an effective tool to embed cultural meanings. Expressing certain concepts in original language can convey subtle cultural and linguistic nuances that can be lost in translation, and knowledge related to certain language are more likely to be more memorized in its own language. As shown in Figure 1, when a human English-Korean bilingual is given a question that is closely related to Korean culture, a question in English and Korean code-switching is more capable of recalling knowledge about “몽유도원도”111A landscape painting by An Gyeon in the early Joseon Dynasty requested by Prince Anpyeong, after his dream about Shangri-la. The painting is drawn on silk with ink., because the concept is more familiar in Korean than in English.
@@
This observation raises intriguing insight about the impact of code-switching in multilingual societies and the potential for equivalent effect in LLMs. Given that code-switching facilitates target language-specific knowledge in human communications, we investigate whether the same applies to English-centric LLMs when solving low-resource language tasks. Therefore, we ask ourselves the following research question: Can code-switched texts activate language-specific knowledge, or turn on a “knowledge switch” in LLMs? By knowledge activation, we refer to the overall process of identifying what knowledge is required, and applying knowledge to answer the question.
@@
To answer the question, we subdivide knowledge activation process into two tasks: (1) In Knowledge Identification task, we investigate if querying LLMs in CS and English yield different knowledge from its encoded memory. Specifically, we evaluate the quality of knowledge from different linguistic settings in terms of faithfulness and helpfulness. (2) In Knowledge Leveraging task, we observe if LLMs can faithfully ground on identified knowledge for solving question-answering (QA) task.
There have been continuous, if not abundant, researches on code-switching in the field of computational linguistics (Aguilar et al., 2020; Rizvi et al., 2021). Recently, after the emergence of LLMs with impressive multilingual abilities, a line of work have discovered LLMs’ abilities in CS (Huzaifah et al., 2024; Yong et al., 2023; Zhang et al., 2023a). However, the focus of such works are only limited to understanding and generating CS of LLMs, while the effectiveness of CS in tasks that involve low-resource language has not yet been explored. To the best of our knowledge, this work is the first to comprehensively analyze the effectiveness of code-switching on knowledge activation to LLMs.
@@
Meanwhile, a crucial challenge when it comes to code-switching is the data scarcity. There is a limited number of CS datasets, let alone culture-focused data (Doğruöz et al., 2021). Since CS often happens in conversations, data are not easily available and the quality is not ensured. To address the shortage of data, efforts have been made to synthetically generate code-switching corpus based on linguistic theories (Pratapa et al., 2018; Rizvi et al., 2021; Salaam et al., 2022). However, these works rely on syntactic parsers and part-of-speech taggers that support limited languages, and the quality of text are highly dependent on the performances of those tools. Therefore, we first construct EnKoQA, a synthetic English-Korean code-switching dataset to explore the potential of CS in low-resource language task. Following Matrix Language Frame Model (Myers-Scotton, 1997), we synthesize Korean QA datasets (Kim et al., 2024b; Son et al., 2024) that encompass various aspects of Korea into English-Korean code-switched questions.
@@
We conduct experiments with EnKoQA and provide extensive analysis on a wide range of multilingual LLMs. The experimental results reveal that CS is able to faithfully activate language-specific knowledge that are encoded in multilingual LLMs compared to high-resource language and target language translation; this tendency was more prominent on domains that specifically requires knowledge in target language and culture.
@@
To the best of our knowledge, this work is the first to comprehensively analyze the effectiveness of code-switching on knowledge activation to LLMs by introducing two tasks.
We propose a qualified English-Korean code-switching QA dataset that is synthesized upon two Korean-centric datasets, and conduct extensive experiments on various families of multilingual LLMs.
Experimental results on extensive LLMs indicate that code-switching has advantages in knowledge activation especially on language-specific domains, suggesting the potential of code-switching text as a tool for conveying cultural nuances in target language tasks.
@@ 2.1 Code-Switching Theories
Many linguistic theories attempt to explain the grammatical construction of code-switched text, such as Equivalence Constraint (EC) theory and Free Morpheme Constraint (FMC) theory proposed by Poplack (1980). EC theory suggests that code-switching occurs at points in a sentence where the structures of both languages are grammatically compatible. FMC theory suggests that code-switching cannot occur between a bound morpheme and a lexical base. (e.g., “He is look-ando for a book.” is a wrong code-switch.)
@@ 2.4 Code-Switched Data Synthesis
3 EnKoQA: English-Korean Code-Switching QA Testset
To compare the effectiveness of code-switching with dominant language and translation in target language when performing language-specific tasks, we introduce EnKoQA, a synthetic English-Korean code-switching dataset that is designed based on MLF model. In this section, we first discuss the details of data construction (§ 3.1), and evaluate performances of LLMs on the dataset (§ 3.2, 3.3).
3.1 Dataset Construction
Data Sources.
We leverage two multiple-choice Korean-centric question-answering datasets that encompass various aspects of Korean language and culture. CLIcK (Kim et al., 2024b) is a Korean benchmark dataset designed to test Korean cultural and linguistic knowledge collected from various official Korean exams and textbooks, e.g., College Scholastic Ability Test of Korea (CSAT). HAE-RAE (Son et al., 2024) is a Korean benchmark dataset originally crafted to capture cultural and contextual nuances inherent to the Korean language, sourced from official Korean exams, textbooks, and text on the internet. In this work, we focus on categories about Korean society to evaluate the effect of CS on activating Korean-specific knowledge. Specifically, we collect 1,995 pairs of eight categories from CLIcK, and 1,027 pairs of five categories from HAE-RAE, resulting in 2,372 QA pairs in nine categories: Popular, Economy, Politics, Tradition, General Knowledge, Society, Geography, History, and Law. More details of original datasets are provided in Appendix A.1.
| GPT-4o | CS | 91.53 | 78.41 | 69.04 | 74.79 | 55.86 | 90.48 | 95.12 | 63.70 | 85.14 | 78.23 |
| EN | 89.83 | 75.00 | 66.19 | 61.97 | 52.64 | 84.52 | 95.12 | 60.40 | 74.32 | 73.33 | |
| KOt | 89.83 | 71.59 | 60.14 | 63.03 | 48.74 | 85.71 | 92.68 | 56.44 | 75.23 | 71.49 | |
@@ 3.1 Dataset Construction
| EN | 76.27 | 46.02 | 49.47 | 38.46 | 42.30 | 69.05 | 73.17 | 52.15 | 63.51 | 56.71 | |
| KOt | 76.27 | 42.05 | 41.99 | 34.62 | 40.23 | 71.43 | 82.93 | 51.98 | 58.11 | 55.51 |
@@
As most LLMs are trained on English-dominant corpora, we regard the English-centric LLM as a bilingual whose matrix language is English but also fairly competent in Korean. To generate code-switched text that follows the MLF model, we need parallel data in Korean and English to extract semantically important words or phrases from Korean text and embed into English text. We first automatically translate all Korean query-choices pairs into English using gpt-3.5-turbo, where the model is instructed to translate the query and choices to English with an one-shot demonstration. Lastly, human supervision was done to ensure translation quality.
@@
Now that we obtain parallel data in both languages, the next step is to embed Korean content morphemes into English sentence. As code-switching mostly happens spontaneously, there does not exist a certain formula for mixing two languages. Moreover, replacing every content word with its Korean equivalent may seem rather artificial. To address this, we simulate a natural code-switching by creating various versions of code-switched texts at different ratios (30, 50, 70, and 90%), then selecting a version that represents the best quality and most naturalness. Specifically, given a question in both languages and a specified proportion, gpt-3.5-turbo identifies content words from the Korean question and integrates them into the English question according to the specified proportion. To collect contexts of various semantic importance, we employ two prompts that define “content word” differently; one defines content words as noun phrases, while the other identifies them as semantically important elements within the context. Eight code-switched candidates are collected per question, from which human annotators select a single candidate that most faithfully follows MLF structure. Comprehensive details about dataset construction are provided in Appendix A.
@@ 3.2 Experimental Settings
Baselines.
To compare performances of LLMs in various language settings, we evaluate on CS, English, and translated Korean (KOt) questions. Korean translation baseline simulates cases where machine translation is adopted to convert task data from English into the target language in low-resource language tasks. We back-translate English translation text to Korean using gpt-3.5-turbo. Prompts that we used for inference are provided in Table 13. We also conduct experiments on the original Korean questions, but do not consider it as major baseline, because we aim to examine the effect of code-switching compared to dominant language, rather than demonstrate the performance of low-resource language. Please refer to Appendix B.3 for further discussions.
@@ 3.3 Results
As shown in Total column from Table 1, the performance on CS significantly outperforms English and KOt across most LLMs in average. The gap between CS and other baselines is especially prominent in GPT-4o and Claude 3.5, where CS peaks in all domains.
@@
While CS outperforms other baselines in many domains, it is worth noting that the gap between CS and English is substantially large on language-sensitive domains such as History and Tradition, both of which target language is essential for preserving information or terminology. Even Llama3 and Gemma2 models which relatively do not perform well on CS questions, show higher scores on CS for such domains. On the other hand, the phenomenon is less consistent for general knowledge (e.g., Society, General), and domains that require expert-level knowledge (e.g., Politics, Law).
CS surpasses translated Korean on most models.
We compare code-switching with translated Korean translation to observe whether CS has advantages in minimizing translation errors. Except for Solar, KOt generally shows lowest performance among three baselines. This suggests that while translating task in target language is not the best practice, CS can faithfully encapsulate meanings and linguistic cues that may be lost in translation, highlighting the potential of leveraging CS for performing non-dominant language tasks.
Ratios do not affect performance.
To ensure that the ratio of code-switching does not influence models’ performances and our dataset is constructed under fair process, we calculate Code-Mixing Index (CMI) scores (Srivastava and Singh, 2021) and report corresponding accuracy in Tradition and History domains. As shown in Table 3, we can see that accuracy scores are quite evenly distributed across all ratios, suggesting that there is no distinct tendency between CMI and accuracy.
@@ 4 Can Code-Switched Questions Activate a “Knowledge Switch” in LLMs?
From Section 3.3, we observe that most LLMs are able to answer correctly to questions in CS than in other baselines. To further investigate on the effectiveness of CS in activating language-specific knowledge, we formulate two tasks: Knowledge Identification and Knowledge Leveraging. We evaluate the tasks in CS and English questions, the two baselines that share the same matrix language.
4.1 Knowledge Identification
Task Description.
When a human English-Korean bilingual is given a question about Korean culture, they will first try to identify what specific knowledge is required to answer the question, and then apply the knowledge to find the correct answer. Depending on which language the question is written in, the quantity and quality of the knowledge may vary, as described in Figure 1. Language-specific knowledge is likely to be encoded much abundantly in its own language, so reading the question in CS will allow more effective knowledge activation than in English. In this sense, knowledge identification task evaluates LLMs’ ability to identify what knowledge is prerequisite for the question. Specifically, the LLM is asked to write a list of factual knowledge that are necessary for solving the given question in one or two sentences.
Evaluation Criteria.
For a qualitative analysis on knowledge identification, we evaluate the quality of a knowledge list based on two criteria: Faithfulness evaluates whether the generated knowledge is factually correct and the model does not output hallucination. Helpfulness evaluates whether the knowledge is relevant to the question, and helpful for answering the question correctly.
4.2 Knowledge Leveraging
Task Description.
@@ 4.2 Knowledge Leveraging
4.3 Experimental Setup